An applied force of 50N is used to accelerate an object with a weight of 80N to the right across a frictional (with friction) surface. The object encounters 10 N of friction. Fill in the blanks below, finding what’s called for.

FN=___

Ff=____

Fapp=___

Fg=_____

m=___

Fnet=___

accerleration= ___


1
Expert's answer
2021-03-14T19:20:12-0400

(a) "F_N=80\ N."

(b) "F_{fr}=10\ N."

(c) "F_{app}=50\ N."

(d) "F_g=80\ N."

(e) "m=\dfrac{F_g}{g}=\dfrac{80\ N}{9.8\ \dfrac{m}{s^2}}=8.16\ kg."

(f) "F_{net}=F_{app}-F_{fr}=50\ N-10\ N=40\ N."

(g) "a=\dfrac{F_{net}}{m}=\dfrac{40\ N}{8.16\ kg}=4.9\ \dfrac{m}{s^2}."